# The shadow and observational images of the non-singular rotating black holes in loop quantum gravity

### v1

@@ 3 Shadows of the rotating LQG black holes
In addition, we can also approximately estimate the angular radius of the rotating black holes, and further present the effect of $\lambda$ on it. In general, the angular radius, $\Theta_{BH}$, is defined as $\Theta_{BH}=\tilde{R}_{s}\frac{\mathcal{M}}{D_{O}}$, where $D_{O}$ is the distance between black hole and observer, and the radius $\tilde{R}_{s}$ with its screen located at the position of black hole is related to the screen’s shadow radius $R_{s}$ obtained in Eq.(30), which can be obtained by using the simple geometrical relationship. As described by [17, 89, 90], when a black hole with mass $\mathcal{M}$ is located far from the observer, the angular radius $\Theta_{BH}$ observed can be quantitatively described as $\Theta_{BH}=9.87098\tilde{R}_{s}\left(\frac{\mathcal{M}}{M_{\astrosun}}\right)\left(\frac{1kpc}{D_{O}}\right)\mu$ $as$. Using this form, we take the $SgrA^{*}$ and $M87^{*}$ as examples, to calculate the angular radius of the black hole in the background of Eq.(9). For the $M87^{*}$, the distance from the Earth is $D_{O}=16.8Mpc$, and the estimated mass of black hole is $\mathcal{M}=(6.5\pm 0.7)\times 10^{6}{M_{\astrosun}}$. And, the the actual shadow diameter should be $\Theta_{M87^{*}}=(37.8\pm 2.7)$$\mu$$as$, which is a $10\%$ discrepancy between the image and shadow diameters [91]. For the $SgrA^{*}$, the observer distance, the estimated mass of black hole, and the shadow diameter are $D_{O}=8kpc$, $\mathcal{M}=(4.0_{-0.6}^{+1.1})\times 10^{6}{M_{\astrosun}}$ and $\Theta_{SgrA^{*}}=(48.7\pm 7)$$\mu$$as$ [92], respectively.
@@
Figure 6: The shadow angular diameter of rotating regular black hole for $a=0.5$.
@@ 4 Shadows illuminated by the celestial light source
In this section, we will employ the backward ray-tracing method to characterize the image of the rotating black hole in the framework of a celestial light source. For the celestial sphere light source model, the black hole is located at the center of the celestial sphere, with its rotation axis pointing towards the North Pole, and is also considerably smaller than both the celestial sphere and the distance between the observer and the origin. The observer is assumed to be situated in the equatorial plane within the sphere. And, the celestial sphere is marked with four into four quadrants, which correspond to four distinct color designations: the green quadrants ($20\leq\theta\leq\pi/$ 2 and $0\leq\phi\leq\pi$), the red quadrants ($20\leq\theta\leq\pi/$ 2 and $\pi\leq\phi\leq 2\pi$), the blue quadrants ($\pi/2\leq\theta\leq\pi$ and $0\leq\phi\leq\pi$), and the yellow quadrants ($\pi/2\leq\theta\leq\pi$ and $\pi\leq\phi\leq 2\pi$). The backward ray-tracing method allows one to trace fewer light rays, without concerning those emitted from the light source that do not reach to the observer. It provides a more convenient way to study black hole images. We follow the numerical strategy outlined in [87], which consists of two parts: the first is a fisheye camera model, primarily used for stereographically projecting the momentum of photons ($p_{\mu}$) onto the screen; the second is the integration of the equations of motion, where the photon’s motion equations are integrated backward along the null geodesics from the observer.
The equations of motion from the Hamiltonian formulation can be expressed as,
@@
where, the overdot implies differentiation with respect to the affine parameter $\tau$. By using the backward ray-tracing method, once the position of the light ray reaching the celestial sphere is determined, its corresponding color is also identified, where the light rays that reach the event horizon being marked in black. In the background of the metric (9), we first solve the above equation (34). By employing the fisheye camera model, then we can successfully obtain the shadow region of the black hole on the screen under the celestial sphere light source, as well as the distortion behavior of light rays around the black hole.
@@
Figure 7: The shadow regions of rotating black hole.
The Fig.1 shows that the maximum values of the non-singular black hole parameter $\lambda$ are approximately $0.923$ and $0.111$ when $a=0.1$ and $a=0.998$, respectively. Based on the parameter constraints, we have presented the black hole shadow regions under the four-color celestial sphere light source, where the symbol $fov$ represents the field of view of the camera, and the observer is positioned at $(r_{o},\theta_{o})$. As shown in Fig.7, when $a=0.1$, the shape of the black hole shadow is essentially circular, and its size decreases with the increase of $\lambda$. When $a=0.998$, the shadow takes on a ”D” shape, and its size also decrease with the increase $\lambda$. It is important to note that, compared to the case when $a=0.1$, the range of variation for parameter $\lambda$ is smaller when $a=0.998$, resulting in a more limited reduction in the shadow area. Additionally, we observe that the impact of parameter $\lambda$ on the shadow region and the surrounding image appears to be quite regular, without any unusual changes such as the cuspy shadow observed in [9].
@@ 5 Images illuminated by the thin accretion disk
In this paper, we assume that the accretion disk starts from the event horizon of the black hole ($r_{h}$) and extends to a very distant point ($r_{f}$), with $r_{f}$ is set to be 10000. And, the position of the observer ($r_{O}$) is located in the region $r_{+}\ll r_{o}<r_{f}$. To obtain the image of the black hole in the background of this thin disk, we first need to find the location of the ISCO. Generally, the position of the ISCO is determined by the following conditions:
@@
$u^{t}_{in}=\left(-g^{tt}\tilde{E}_{ISCO}+g^{t\phi}\tilde{L}_{ISCO}\right)_{\mid_{\theta=\pi/2}},\quad$ $u^{\phi}_{in}=\left(-g^{t\phi}\tilde{E}_{ISCO}+g^{\phi\phi}\tilde{L}_{ISCO}\right)_{\mid_{\theta=\pi/2}},$ (42)
$u^{r}_{in}=-\sqrt{-\frac{g_{tt}u^{t}_{in}u^{t}_{in}+2g_{t\phi}u^{t}_{in}u^{\phi}_{in}+g_{\phi\phi}u^{\phi}_{in}u^{\phi}_{in}+1}{g_{rr}}}_{\mid_{\theta=\pi/2}},$ $u^{\theta}_{in}=0.$ (43)
@@
As previously mentioned, when we trace the light rays backward from the observer’s position, the rays may intersect the equatorial plane’s accretion disk once ($n=1$), twice ($n=2$), or even more times ($n>2$). Each intersection allows the observer to receive additional luminosity, with the intersection position denoted as $r_{n}$, which we define as the transfer function. Thus, by neglecting reflection effects and the thickness of the accretion disk, the intensity observed on the observer’s screen can be determined as follows:
@@
where $N$ is the maximum number of intersections between the light rays and the accretion disk, and $f_{n}$ is the fudge factor which is fixed to 1 in this paper . And, by considering that the observational wavelength of the black hole images taken by the EHT is $1.3mm$ ($230GHz$), we choose the emissivity of the thin disk as a second-order polynomial in log-space,
$J(r)=exp\left[{-\frac{1}{2}z^{2}-2z}\right],\quad z=log\frac{r}{r_{+}}.$ (45)
On the other hand, the redshift factor is $g_{n}=\nu_{o}/\nu_{n}$ with $\nu_{o}$ is the observed frequence on the screen, and $\nu_{n}$ denotes the frequence observed by the local rest frames comoving with the accretion disk. Since the accretion disk is an electrically neutral plasma and moves along timelike geodesics with $\tilde{E}$ and $\tilde{L}$, the redshift factor of these circularly orbiting fluids outside the ISCO can be written as follows,
@@
(e) (a): The position of observer $\theta_{obs}=80^{\circ}$
@@
(e) (b): The position of observer $\theta_{obs}=163^{\circ}$
@@
(e) (a): The redshifts of direct images of the accretion disk for $\theta_{obs}=80^{\circ}$
@@
(e) (b): The redshifts of lensed images of accretion disk for $\theta_{obs}=80^{\circ}$.
@@
(e) (c): The redshifts of direct images of accretion disk $\theta_{obs}=163^{\circ}$.
@@
(e) (d): The redshifts of lensed images of accretion disk $\theta_{obs}=163^{\circ}$.
@@
(e) (a): The position of observer $\theta_{obs}=80^{\circ}$
@@
(e) (b): The position of observer $\theta_{obs}=163^{\circ}$
@@
At the position $80^{\circ}$, it shows in Fig.12 that there is a noticeable asymmetry in the intensity distribution along the X-direction on the screen, while the Y-direction remains relatively symmetric. The rotation parameter $a$ causes some variation in the peak intensity along the X-direction, with corresponding differences in the X-coordinates. More significantly, the quantum parameter has a pronounced impact on the intensity distribution. Whether along the X-axis or Y-axis, and whether $a=0.1$ or $a=0.998$, an increase in $\lambda$ leads to a reduction in both the intensity and its peak value, along with a decrease in the spacing of the coordinates corresponding to the peaks. Moreover, this also holds true for the case at $163^{\circ}$. Combining the analysis of the black hole shadow size in Section 3, the discussion of celestial images in Section 4, and the study of thin disk imaging in this section, we can conclude that, compared to Kerr black holes, the quantum parameter $\lambda$ of non-singular rotating black holes in LQG weakens the strength of gravitational field. This results in a reduction of the black hole shadow, a smaller inner shadow, and a reduction in the intensity of the observable thin disk image.
For the retrograde case, we present the accretion disk images, redshift images, and intensity distribution on the screen by using $a=0.998,\lambda=0.01$ and $a=0.998,\lambda=0.1$ as examples.
@@
(e) (a): The images and intensity for $\theta_{obs}=80^{\circ}$ and $a=0.998$
@@
(c) (b): The position of observer $\theta_{obs}=80^{\circ}$
@@
(d) (a): The images for $\theta_{obs}=80^{\circ}$ and $a=0.94$
@@
(d) (b): The redshift of derect images for $\theta_{obs}=80^{\circ}$ and $a=0.94$
@@
(c) (b): The intensity of images for $\theta_{obs}=80^{\circ}$ and $a=0.94$
@@ 6 Conclusions and discussions
In this paper, based on the ray-tracing method, we investigate the shadows of non-singular rotating black holes in LQG, and further conduct numerical simulations of the inner shadows, celestial sphere images, and thin disk images of these black holes. The non-singular rotating spacetimes encompasses three scenarios: (A) the rotating wormhole without the event horizon; (B) the black hole with only an outer event horizon; and (C) the rotating black hole with both an outer and inner event horizons. Since the astronomical observations have ruled out the possibility of a rotating wormhole, we focus on scenarios (B) and (C) for our study. Firstly, we studied the shadow of non-singular rotating black holes using stereographic projection techniques in the ZAMO frame, and constrained the relevant parameters of these black holes using astronomical observation datas from M87* and SgrA*. Then, by considering a celestial sphere light source model, we continue to investigate the images of celestial sphere light sources and the behavior of light distortion near the black hole. Finally, by assuming the presence of an optically thin geometrical accretion disk in the equatorial plane of black hole, we also studied the characteristics of the inner shadow, thin disk images, and redshift profiles of black hole under both the prograde and retrograde disks.
@@
When the thin accretion disk is located at the equatorial plane with the prograde angular velocity, for an observer with $\theta_{obs}=80\pi/180$, the results show that the spin parameter $a$ plays a dominant role in shaping the inner shadow of black hole, and its influence on the size of the inner shadow is relatively weak. On the other hand, the quantum parameter $\lambda$ has little effect on the shape of black hole shadow but plays a dominant role in determining the size of the inner shadow. Moreover, for the image of non-singular rotating black holes, the direct image and the lensed image are distinguishable. When the observer is located at $163^{\circ}$, the inner shadow of black hole appears almost as a black disk, and at this point, the influence of $a$ and $\lambda$ on the black hole’s shadow is reflected only in its size. Comparatively, the effect of $\lambda$ is more pronounced. At the same time, the direct image and the lensed image are nearly indistinguishable at this inclination. For the observed intensity, whether at $80^{\circ}$ or $163^{\circ}$, the intensity along the Y-axis is nearly symmetrical, while along the X-axis, it is stronger on the left and weaker on the right. Both $a$ and $\lambda$ cause the two peaks of the intensity to move closer together, though the influence of $\lambda$ is more pronounced. Furthermore, $\lambda$ has a more noticeable effect on the observed intensity, primarily reducing the intensity distribution area and its maximum value. Additionally, it further shows that both $a$ and $\lambda$ mainly decrease the size of the shadow of redshift images, and reduce the area of the redshift images. Furthermore, based on the analysis of retrograde accretion disk images and redshift images of non-singular rotating black holes, it is found that the influence of $a$ and $\lambda$ on the inner shadow, observed intensity, and redshift images remains consistent with that in the prograde case. Finally, we attempted to examine whether the appearance of thin disk accretion for non-singular rotating black holes can reflect their different internal structures. When fixing $a=0.94$, $\lambda=0.0539$ corresponds to case (B), $\lambda=0.0537$ corresponds to case (C), and $\lambda=0.0538$ is at the critical value. It turns out that the black hole’s accretion disk images, redshift images, and intensity distribution curves on the screen in all three cases reveal no discernible differences. This indicates that the thin disk accretion images of black hole cannot effectively reflect the internal structure of black hole.

### v2

@@ 3 Shadows of the rotating LQG black holes
In addition, we can also approximately estimate the angular radius of the rotating black holes, and further present the effect of $\lambda$ on it. In general, the angular radius, $\Theta_{BH}$, is defined as $\Theta_{BH}=\tilde{R}_{s}\frac{\mathcal{M}}{D_{O}}$, where $D_{O}$ is the distance between black hole and observer, and the radius $\tilde{R}_{s}$ with its screen located at the position of black hole is related to the screen’s shadow radius $R_{s}$ obtained in Eq.(29), which can be obtained by using the simple geometrical relationship. As described by [18, 112, 113], when a black hole with mass $\mathcal{M}$ is located far from the observer, the angular radius $\Theta_{BH}$ observed can be quantitatively described as $\Theta_{BH}=9.87098\tilde{R}_{s}\left(\frac{\mathcal{M}}{M_{\astrosun}}\right)\left(\frac{1kpc}{D_{O}}\right)\mu$$as$. Using this form, we take the SgrA* and M87* as examples, to calculate the angular radius of the black hole in the background of Eq.(8). For the $M87^{*}$, the distance from the Earth is $D_{O}=16.8Mpc$, and the estimated mass of black hole is $\mathcal{M}=(6.5\pm 0.7)\times 10^{6}{M_{\astrosun}}$. And, the the actual shadow diameter should be $\Theta_{M87^{*}}=(37.8\pm 2.7)$$\mu$$as$, which is a $10\%$ discrepancy between the image and shadow diameters [114]. For the $SgrA^{*}$, the observer distance, the estimated mass of black hole, and the shadow diameter are $D_{O}=8kpc$, $\mathcal{M}=(4.0_{-0.6}^{+1.1})\times 10^{6}{M_{\astrosun}}$ and $\Theta_{SgrA^{*}}=(48.7\pm 7)$$\mu$$as$ [115], respectively.
@@
Figure 6: The shadow angular diameter of non-singular rotating black hole for $a=0.5$.
@@ 4 Shadows illuminated by the celestial light source
In this section, we will employ the backward ray-tracing method to characterize the image of the rotating black hole in the framework of a celestial light source. For the celestial sphere light source model, the black hole is located at the center of the celestial sphere, with its rotation axis pointing towards the North Pole, and is also considerably smaller than both the celestial sphere and the distance between the observer and the origin. The observer is assumed to be situated in the equatorial plane within the sphere. And, the celestial sphere is marked with four into four quadrants, which correspond to four distinct color designations: the green quadrants ($20\leq\theta\leq\pi/$ 2 and $0\leq\phi\leq\pi$), the red quadrants ($20\leq\theta\leq\pi/$ 2 and $\pi\leq\phi\leq 2\pi$), the blue quadrants ($\pi/2\leq\theta\leq\pi$ and $0\leq\phi\leq\pi$), and the yellow quadrants ($\pi/2\leq\theta\leq\pi$ and $\pi\leq\phi\leq 2\pi$). The backward ray-tracing method allows one to trace fewer light rays, without concerning those emitted from the light source that do not reach to the observer. It provides a more convenient way to study black hole images. We follow the numerical strategy outlined in [110], which consists of two parts: the first is a fisheye camera model, primarily used for stereographically projecting the momentum of photons $p_{\mu}$ onto the screen; the second is the integration of the equations of motion, where the photon’s motion equations are integrated backward along the null geodesics from the observer.
The equations of motion from the Hamiltonian formulation can be expressed as
@@
where the overdot implies differentiation with respect to the affine parameter $\tau$. By using the backward ray-tracing method, once the position of the light ray reaching the celestial sphere is determined, its corresponding color is also identified, where the light rays that reach the event horizon being marked in black. In the background of the metric (8), we first solve the above equation (33). By employing the fisheye camera model, then we can successfully obtain the shadow region of the black hole on the screen under the celestial sphere light source, as well as the distortion behavior of light rays around the black hole.
@@
Figure 7: The shadow regions of non-singular rotating black hole.
The Fig.1 shows that the maximum values of the non-singular black hole parameter $\lambda$ are approximately $0.923$ and $0.111$ when $a=0.1$ and $a=0.998$, respectively. Based on the parameter constraints, we have presented the black hole shadow regions under the four-color celestial sphere light source, where the symbol $fov$ represents the field of view of the camera, and the observer is positioned at $(r_{o},\theta_{o})$. As shown in Fig.7, when $a=0.1$, the shape of the black hole shadow is essentially circular, and its size decreases with the increase of $\lambda$. When $a=0.998$, the shadow takes on a “D” shape, and its size also decrease with the increase $\lambda$. It is important to note that, compared to the case when $a=0.1$, the range of variation for parameter $\lambda$ is smaller when $a=0.998$, resulting in a more limited reduction in the shadow area. Additionally, we observe that the impact of parameter $\lambda$ on the shadow region and the surrounding image appears to be quite regular, without any unusual changes such as the cuspy shadow observed in [9].
@@ 5 Images illuminated by the thin accretion disk
In this paper, we assume that the accretion disk starts from the event horizon of the black hole ($r_{h}$) and extends to a very distant point ($r_{f}$), with $r_{f}$ is set to be 10000. And, the position of the observer ($r_{O}$) is located in the region $r_{+}\ll r_{o}<r_{f}$. To obtain the image of the black hole in the background of this thin disk, we first need to find the location of the ISCO. Generally, the position of the ISCO is determined by the following conditions
@@
$u^{t}_{in}=\left(-g^{tt}\tilde{E}_{ISCO}+g^{t\phi}\tilde{L}_{ISCO}\right)_{\mid_{\theta=\pi/2}},$
$u^{r}_{in}=-\sqrt{-\frac{g_{tt}u^{t}_{in}u^{t}_{in}+2g_{t\phi}u^{t}_{in}u^{\phi}_{in}+g_{\phi\phi}u^{\phi}_{in}u^{\phi}_{in}+1}{g_{rr}}}_{\mid_{\theta=\pi/2}},$
$u^{\theta}_{in}=0,$
$u^{\phi}_{in}=\left(-g^{t\phi}\tilde{E}_{ISCO}+g^{\phi\phi}\tilde{L}_{ISCO}\right)_{\mid_{\theta=\pi/2}}.$ (41)
@@
As previously mentioned, when we trace the light rays backward from the observer’s position, the rays may intersect the equatorial plane’s accretion disk once ($n=1$), twice ($n=2$), or even more times ($n>2$). Each intersection allows the observer to receive additional luminosity, with the intersection position denoted as $r_{n}$, which we define as the transfer function. Thus, by neglecting reflection effects and the thickness of the accretion disk, the intensity observed on the observer’s screen can be determined as follows
@@
where $N$ is the maximum number of intersections between the light rays and the accretion disk, and $f_{n}$ is the fudge factor which is fixed to 1 in this paper . And, by considering that the observational wavelength of the black hole images taken by the EHT is $1.3mm$ ($230GHz$), we choose the emissivity of the thin disk as a second-order polynomial in log-space
$J(r)=exp\left[{-\frac{1}{2}z^{2}-2z}\right],\quad z=\log\frac{r}{r_{+}}.$ (43)
On the other hand, the redshift factor is $g_{n}=\nu_{o}/\nu_{n}$ with $\nu_{o}$ is the observed frequence on the screen, and $\nu_{n}$ denotes the frequence observed by the local rest frames comoving with the accretion disk. Since the accretion disk is an electrically neutral plasma and moves along timelike geodesics with $\tilde{E}$ and $\tilde{L}$, the redshift factor of these circularly orbiting fluids outside the ISCO can be written as follows
@@
(e) (a) The position of observer $\theta_{obs}=80^{\circ}$.
@@
(e) (b) The position of observer $\theta_{obs}=163^{\circ}$.
@@
(e) (a) The redshifts of direct images of the accretion disk for $\theta_{obs}=80^{\circ}$.
@@
(e) (b) The redshifts of lensed images of accretion disk for $\theta_{obs}=80^{\circ}$.
@@
(e) (c) The redshifts of direct images of accretion disk $\theta_{obs}=163^{\circ}$.
@@
(e) (d) The redshifts of lensed images of accretion disk $\theta_{obs}=163^{\circ}$.
@@
(e) (a) The position of observer $\theta_{obs}=80^{\circ}.$
@@
(e) (b) The position of observer $\theta_{obs}=163^{\circ}$
@@
At the position $80^{\circ}$, it shows in Fig.12 that there is a noticeable asymmetry in the intensity distribution along the X-axis on the screen, while the Y-axis remains relatively symmetric. The rotation parameter $a$ causes some variation in the peak intensity along the X-axis, with corresponding differences in the X-coordinates. More significantly, the quantum parameter has a pronounced impact on the intensity distribution. Whether along the X-axis or Y-axis, and whether $a=0.1$ or $a=0.998$, an increase in $\lambda$ leads to a reduction in both the intensity and its peak value, along with a decrease in the spacing of the coordinates corresponding to the peaks. Moreover, this also holds true for the case at $163^{\circ}$. Combining with the above analysis, we can conclude that, compared to Kerr black holes, the quantum parameter $\lambda$ of non-singular rotating black holes in LQG weakens the strength of gravitational field. This results in a reduction of the black hole shadow, a smaller inner shadow, and a reduction in the intensity of the observable thin disk image.
For the retrograde case, we present the accretion disk images, redshift images, and intensity distribution on the screen by using ($a=0.998,\lambda=0.01$) and ($a=0.998,\lambda=0.1$) as examples.
@@
(e) (a) The images and intensity for $\theta_{obs}=80^{\circ}$ and $a=0.998$.
@@
(c) (b) The position of observer $\theta_{obs}=80^{\circ}$
@@
(d) (a) The images for $\theta_{obs}=80^{\circ}$ and $a=0.94.$
@@
(d) (b) The redshift of derect images for $\theta_{obs}=80^{\circ}$ and $a=0.94$
@@
(c) (b) The intensity of images for $\theta_{obs}=80^{\circ}$ and $a=0.94$.
@@ 6 Conclusions and discussions
In this paper, based on the ray-tracing method, we investigate the shadows of non-singular rotating black holes in LQG, and further conduct numerical simulations of the inner shadows, celestial sphere images, and thin disk images of these black holes. The non-singular rotating spacetimes encompasses three scenarios: (Region A) the rotating wormhole without the event horizon; (Region B) the black hole with only an outer event horizon; and (Region C) the rotating black hole with both an outer and inner event horizons. Since the astronomical observations have ruled out the possibility of a rotating wormhole, we focus on scenarios (Region B) and (Region C) for our study. Firstly, we studied the shadow of non-singular rotating black holes using stereographic projection techniques in the ZAMO frame, and constrained the relevant parameters of these black holes using astronomical observation datas from M87* and SgrA*. Then, by considering a celestial sphere light source model, we continue to investigate the images of celestial sphere light sources and the behavior of light distortion near the black hole. Finally, by assuming the presence of an optically thin geometrical accretion disk in the equatorial plane of black hole, we also studied the characteristics of the inner shadow, thin disk images, and redshift profiles of black hole under both the prograde and retrograde disks.
@@
When the thin accretion disk is located at the equatorial plane with the prograde angular velocity, for an observer with $\theta_{obs}=80\pi/180$, the results show that the spin parameter $a$ plays a dominant role in shaping the inner shadow of black hole, and its influence on the size of the inner shadow is relatively weak. On the other hand, the quantum parameter $\lambda$ has little effect on the shape of black hole shadow but plays a dominant role in determining the size of the inner shadow. Moreover, for the image of non-singular rotating black holes, the direct image and the lensed image are distinguishable. When the observer is located at $163^{\circ}$, the inner shadow of black hole appears almost as a black disk, and at this point, the influence of $a$ and $\lambda$ on the black hole’s shadow is reflected only in its size. Comparatively, the effect of $\lambda$ is more pronounced. At the same time, the direct image and the lensed image are nearly indistinguishable at this inclination. For the observed intensity, whether at $80^{\circ}$ or $163^{\circ}$, the intensity along the Y-axis is nearly symmetrical, while along the X-axis, it is stronger on the left and weaker on the right. Both $a$ and $\lambda$ cause the two peaks of the intensity to move closer together, though the influence of $\lambda$ is more pronounced. Furthermore, $\lambda$ has a more noticeable effect on the observed intensity, primarily reducing the intensity distribution area and its maximum value. Additionally, it further shows that both $a$ and $\lambda$ mainly decrease the size of the shadow of redshift images, and reduce the area of the redshift images. Furthermore, based on the analysis of retrograde accretion disk images and redshift images of non-singular rotating black holes, it is found that the influence of $a$ and $\lambda$ on the inner shadow, observed intensity, and redshift images remains consistent with that in the prograde case. Finally, we attempted to examine whether the appearance of thin disk accretion for non-singular rotating black holes can reflect their different internal structures. When fixing $a=0.94$, $\lambda=0.0539$ corresponds to case (Region B), $\lambda=0.0537$ corresponds to case (Region C), and $\lambda=0.0538$ is at the critical value. It turns out that the black hole’s accretion disk images, redshift images, and intensity distribution curves on the screen in all three cases reveal no discernible differences. This indicates that the thin disk accretion images of black hole cannot effectively reflect the internal structure of black hole.